OpenStudy (anonymous):

Which expression is equivalent to one over fourn − 16? HELP PLEASE FOR A MEDAL! A)1/4(n-4) B)1/4(n-64) C)4(n-4) D)4(n-64)

Nnesha (nnesha):

distribute parentheses by outside term

OpenStudy (anonymous):

Can you tell me the answer I will give you a medal? @Nnesha

Nnesha (nnesha):

distribute = multiply example a(b+c)= a times b + a times c = ab+ac now try that

OpenStudy (anonymous):

Is B the answer? @Nnesha

Nnesha (nnesha):

how ?? 4(n-64) distribute parentheses by outside term let me know what you get :-)

OpenStudy (anonymous):

D @Nnesha

Nnesha (nnesha):

how ??

OpenStudy (anonymous):

I don't get this it's between A and C but how do you do this? @Nnesha

Nnesha (nnesha):

distribute 2nd parentheses by outside term for example B  multiply 4 times n= ?? 4 times -64 =?